OpenStudy (anonymous):

Factor completely: x2 – 64 (x – 64)(x + 1) (x – 64)(x – 1) (x – 8)(x – 8) (x – 8)(x + 8)

OpenStudy (lgbasallote):

this is a difference of two squares

Parth (parthkohli):

Ahhh! Difference of squares: \( \color{Black}{\Rightarrow a^2 - b^2 = ( a + b)( a - b) }\)

OpenStudy (lgbasallote):

note that in difference of two squares \[a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (studyhard):

(x-8)(x+8)

Parth (parthkohli):

yes

OpenStudy (anonymous):

Factor completely: 4x2 – 121 (2x – 11)(2x + 11) (2x – 11)(2x – 11) (4x – 121)(x + 1 ) (4x – 121)(x – 1)